# A Systematic Survey on Large Language Models for Algorithm Design

### v1

@@ Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and superior solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has progressed significantly, finding applications in diverse areas such as optimization, machine learning, mathematical reasoning, and scientific exploration. Given the rapid development and broadening scope of this field, a systematic review is both timely and essential. This paper provides a systematic review of the works on LLM4AD. First, we present an overview and summary of existing studies. Then, we present a systematic categorization, and a review of existing works along four dimensions including the role of LLMs, search techniques, prompt strategies, and application fields. We also discuss the achievements and challenges in each area and the capabilities of LLM4AD in addressing them. Finally, we explore current limitations and propose several open questions and promising directions for future research.
Algorithm design, Large language model, Learn to optimize, Optimization, Heuristic, Hyperheuristic, Evolutionary computation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXcopyright: none
@@ 1. Introduction
Algorithms play a crucial role in addressing a myriad of problems across diverse fields, including industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). The traditional hand-crafted approach to algorithm design requires extensive expert knowledge and considerable effort, which is tedious and time-consuming. There has been a growing interest in the integration of learning and computational intelligence techniques to streamline and enhance the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
@@
In the past three years, the integration of large language models for algorithm design (LLM4AD) has become an emerging area of research that promises to revolutionize how algorithms are conceived, optimized, and implemented. The impressive power and adaptability of LLMs have shown promising in improving and transforming the algorithm design process, such as heuristic generation (Liu et al., 2024d), code optimization (Ma et al., 2024a), and even the creation of novel algorithms (Girotra et al., 2023) tailored to specific problems. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the solutions produced (Liu et al., 2024d; Romera-Paredes et al., 2024).
@@
While LLM4AD is gaining traction, systematic reviews in this nascent field remain sparse. Existing literature primarily focuses on isolated applications of LLMs in specific algorithmic contexts, lacking a systematic overview of the methodologies, applications, challenges, and future directions. Several papers have been carried out to survey the LLM for/with specific optimization topic (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]) and others review the general LLM applications (Hadi et al., 2024) or the application on specific problems including electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), soft engineering (Hou et al., 2023), recommendation (Wu et al., 2024b) and agent (Wang et al., 2024c). This paper aims to fill this gap by providing a systematic review of the current state of large language models in algorithm design. We will explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs in enhancing algorithm design and lays the groundwork for further innovations in this exciting field. The contributions of this paper are as follows:
  • Systematic Review of LLM4AD: We present a systematic review of the developments in using large language models for algorithm design, covering a significant corpus of 180+ research papers published in the last three years. We not only synthesize the current state of research but also categorize the research works, providing a critical analysis of methodologies, results, and algorithm design advancements. We expect it to be a helpful resource for both new researchers entering the field and experienced experts looking for a consolidated update on recent developments.
@@
  • Development of a Multi-dimensional Taxonomy: We introduce a multi-dimensional taxonomy that categorizes the works and functionalities of LLM4AD into four distinct dimensions: 1) Roles of LLMs in algorithm design, which delineates how these models contribute to or enhance algorithm design, including LLMs as optimizers, LLMs as predictors, LLMs as extractors, and LLMs as designers; 2) Search methods, which explores the various approaches used by LLMs to navigate and optimize search spaces in algorithm design; 3) Prompt methods, which examines how diverse prompting strategies are used; and 4) Application domains, which identifies the key fields and industries where LLMs are being applied to solve complex algorithmic challenges. This taxonomy not only clarifies the landscape but also aids in identifying gaps and opportunities for future research.
@@
  • Discussion on Limitations and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. We discuss challenges such as scalability, interpretability, high cost, and security. Furthermore, we propose potential future research directions that could address these limitations, including developing domain-specific LLMs, exploring multi-modal LLMs, implementing interactive systems with human experts, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automatic algorithm design, and establishing benchmarking standards for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
@@ 2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We further delineate the scope of our survey as follows:
  • We do not intend to cover all literature on LLMs and algorithms. Specifically, we exclude works focused on other branches, such as algorithms for optimizing LLMs, including prompt engineering algorithms (Sahoo et al., 2024) and algorithms for training LLMs (Ackerman et al., 2024).
  • The term Large Language Models refers to language models of sufficient scale to enable robust zero-shot performance across various tasks, including linguistic comprehension, code generation, and mathematical reasoning. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023).
    • Studies employing smaller models for algorithm design, such as those prevalent in conventional model-based algorithms and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. While it is challenging to precisely define “large” models, most leading-edge LLMs contain over one billion parameters (Zhao et al., 2023; Minaee et al., 2024).
    • Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
@@ 2.2. Statistics
  • Stage I Data extraction and collection: We collect the related papers through Google Scholar, Web of Science, and Scopus. The logic of our search is the title must include any combinations of at least one of the following two groups of words “LLM”, “LLMs”, “Large Language Model”, “Large Language Models” and “Algorithm”, “Heuristic”, “Search”, “Optimization”, “Optimizer”, “Design”, “Function” (e.g., LLM and optimization, LLMs and algorithm). As a fast-growing research field, the majority of papers on LLM4AD are published online as pre-print manuscripts such as in Arxiv, which results in a notably more paper collected from Google Scholar (800+) than Web of Science and Scopus. After removing duplicated papers, we ended up with 850 papers as of July 1, 2024.
  • Stage II Abstract Scanning: We first check the title and abstract of each paper to efficiently exclude papers that are not relative. The criteria used for exclusion include these papers that are not in English, not for algorithm design and not using large language models. After scanning, 260 papers are remaining.
  • Stage III Full Scanning: In this stage, we carefully check the entire manuscript of each paper. We remove these papers that do not include relevant content in their main paper. After scanning, there are 160 papers left.
  • Stage IV Supplementation: In order to avoid missing any important contributions, we searched the related works manually according to our past knowledge in this field. After integrating the additional papers, we got 180+ papers in the end.
@@
Figure 3. The word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It features the top 80 keywords, organized into four color-coded clusters. The Blue Cluster, which prominently features the keyword “Language”, includes other frequently used terms such as “strategy”, “reasoning”, “prompt”, and “function”, which are crucial for task resolution. The Red Cluster is centered around “GPT”. This cluster also contains terms like “fine-tuning”, “text”, and “accuracy”, which are crucial in model training and inference. The Green Cluster focuses on search and optimization, encompassing terms such as “evolutionary algorithm”, “combination”, and “diversity”. Finally, the Yellow Cluster emphasizes the role of LLMs in scientific discovery, highlighting keywords such as “scientist”, “hypothesis”, and “concept”.
We will first present an overview of the LLM4AD paper list and then present a taxonomy to systematically review the progress. In addition to the organized list of papers, this survey also incorporates some publications released after July 1, 2024. We will update the statistics as needed.
@@ 2.3. Overview
Fig. 2(a) illustrates the trend in the number of papers published over time, with the timeline expressed in months. The graph shows a marked rise in research activity related to LLM4AD, particularly noting that most of the studies have been conducted in the last year. This suggests that LLM4AD is an emerging field, and we expect a significant increase in research output in the near future as scholars from diverse fields become aware of its considerable potential. It is worth noting that as a rapidly expanding field, the majority of works are initially published as preprints on Arxiv. Many of these have subsequently been accepted by top conferences such as ICML and NeurIPS.
@@ 2.4. Taxonomy
Research in the field of LLM4AD primarily explores three fundamental components: 1) LLMs, 2) algorithms, and 3) applications. Studies within LLM4AD are categorized based on the variations and interactions of these elements. This paper presents a taxonomy organized into four dimensions: 1) LLM Roles, 2) Search Methods, 3) Prompt Methods, and 4) Applications.
@@
  • Search Methods: This category involves the search methods used. It includes a variety of search strategies, from the basic and widely utilized random search methods to more complex approaches like single-point-based, population-based, and uncertainty-guided methods.
@@ 3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an optimization algorithm to generate and refine solutions (Fig. 5). This innovative use of LLMs extends beyond their traditional applications in natural language processing to more complex problem-solving tasks that involve searching and optimization. The process typically involves employing the LLM in an iterative search framework, where the model actively participates in proposing potential solutions based on feedback received at each step. The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions, which might be infeasible with traditional optimization methods alone.
@@
One of the initial efforts to utilize LLMs as optimizers in algorithm design is by Yang et al. (2023). They leverage the in-context learning capabilities of LLMs to generate novel solutions for specific problems, based on previously evaluated solutions. The process of generating solutions is considered a form of in-context learning, where evaluated solutions serve as contextual information. Typically, these solutions are organized to help the LLM understand preferences for optimized outcomes, thereby enabling it to propose superior alternatives. This method is applied iteratively to refine solutions further. Yang et al. (2023) have successfully demonstrated this technique across various domains, including continuous and combinatorial optimization, as well as machine learning tasks.
@@
From an Evolutionary Computation (EC) perspective, using LLMs to generate solutions from existing data can be seen as analogous to search operators in EC algorithms. For instance, Liu et al. (2023a) introduce the use of LLMs as evolutionary operators to tackle multi-objective problems. This method involves breaking down a multi-objective problem into simpler single-objective tasks, with LLMs acting as black-box search operators for each sub-problem, thereby suggesting new solutions. This decomposition framework helps prevent LLMs from being overwhelmed by the complexity of direct multi-objective optimization. In a related study, Liu et al. (2024b) explore the integration of LLMs within evolutionary algorithms (EAs), not just for generating solutions but also for guiding selection, crossover, and mutation processes. Meanwhile, Brahmachary et al. (2024) propose a novel population-based evolutionary framework that includes both exploration and exploitation pools, with solutions being exchanged during the optimization process and LLMs generating solutions for both pools. Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their “EvoLLM” approach, which showed superior performance over baseline algorithms in synthetic BBOB functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms could further improve the performance of EvoLLM. LLMaO has also been used for hyperparameter optimization by Custode et al. (2024), who present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
@@
Beyond traditional optimization tasks, LLMs are increasingly being used in prompt engineering to enhance zero-shot performance, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models. Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity, and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors, as explored in studies by Wang et al. (2023b) and Guo et al. (2023). Reflection-based methods, in particular, mimic the trial-and-error process of skilled human prompt designers. Despite the progress, challenges remain, as Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. Moreover, even semantically correct reflections do not always result in effective prompts in a single refinement step, often due to unpredictable model responses. To address these issues, they proposed a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models. A recent development in this field is the gradient-based prompt optimization method, termed GPO, introduced by Tang et al. (2024c). This method identifies two critical components of gradient-based search: the update direction and the update method. At each step, GPO retrieves relevant prompts from the optimization trajectory to determine the update direction, then applies a generation-based refinement strategy for the update, while managing the edit distance through a cosine-based decay strategy.
@@ 3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models, which serve as stand-ins for more complex or computationally expensive models. As illustrated in Fig. 6, these LLMs are utilized to predict the outcomes or responses, functioning either in a classification or regression context (Hao et al., 2024). In classification tasks, the LLMaP approach involves the LLM predicting categorical labels or classes based on input data. On the other hand, in regression tasks, the LLM predicts continuous outputs rather than categories. Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses based on the training they have received on vast datasets. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies or subtleties in the data (Egami et al., 2024). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
@@
The majority of LLMaP works use LLMs as regression models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024e) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023a) introduce CodonBERT, LLMs for mRNAs, which is used for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and was trained on over 10 million mRNA sequences from various organisms. The model can be used for predicting mRNA properties and outperforms previous methods, as demonstrated on a flu vaccine dataset. Another application on battery electrolytes design by Soares et al. (2024) demonstrates the use of transformer-based models in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a)
Several studies have explored the application of LLMs in classification tasks. For instance, Hao et al. (2024) introduce the LAEA model, which employs LLMs as surrogate models within evolutionary algorithms. This approach is used for both classification, determining the quality of solutions as either good or bad, and regression, by approximating the value of solutions. This method significantly aids in solving optimization problems by evaluating new solutions based on historical data, thus eliminating the need for costly model training. In another study, Chen et al. (2023b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance.
@@
Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in reinforcement learning (RL). Their method demonstrated effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.
Other works study LLM-based classification models. For example, Hao et al. (2024) introduce the LAEA, which utilizes LLMs as surrogate models in evolutionary algorithms and investigates both classification (predicting whether a solution is good or bad), and regression (approximating its value). It helps solve optimization problems by evaluating the quality of new solutions based on historical data eliminating expensive model training. Chen et al. (2023b) discuss a label-free node classification method using LLMs to annotate nodes, which are then used to train graph neural networks for enhanced performance. Except for the binary classification task, Bhambri et al. (2024) use LLMs to predict discrete actions for constructing reward shaping functions in RL, demonstrating effectiveness in the BabyAI environment. Wang et al. (2024d) explore the use of LLMs in federated search, employing them in a zero-shot setting to select resources effectively. Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product search, achieving significant improvements by finetuning LLMs.

### v2

@@ Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has progressed significantly, finding applications in diverse areas such as optimization, machine learning, mathematical reasoning, and scientific discovery. Given the rapid development and broadening scope of this field, a systematic review is both timely and essential. This paper provides a systematic review of the works on LLM4AD. First, we present an overview and summary of existing studies. Then, we present a systematic taxonomy and a review of existing works along four dimensions, including the role of LLMs, search techniques, prompt strategies, and applications, with a discussion of the potential and achievements of using LLMs. Finally, we explore current challenges and propose several open questions and promising directions for future research.
Algorithm design, Large language model, Optimization, Heuristic, Hyperheuristic, Evolutionary algorithm.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXcopyright: none
@@ 1. Introduction
Algorithms play a crucial role in addressing various problems across diverse fields, including industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). The traditional hand-crafted approach to algorithm design requires extensive expert knowledge and considerable effort, which is tedious and time-consuming. There has been a growing interest in the integration of learning and computational intelligence techniques to streamline and enhance the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
@@
In the past three years, the integration of Large Language Models for Algorithm Design (LLM4AD) has become an emerging area of research that promises to revolutionize how algorithms are conceived, optimized, and implemented. The impressive power and adaptability of LLMs have shown promising in improving and transforming the algorithm design process, such as heuristic generation (Liu et al., 2024d), code optimization (Ma et al., 2024a), and even the creation of novel algorithms (Girotra et al., 2023) tailored to specific problems. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the produced solutions (Liu et al., 2024d; Romera-Paredes et al., 2024).
@@
While LLM4AD is gaining traction, systematic reviews in this nascent field remain sparse. Existing literature primarily focuses on applications of LLMs in specific algorithmic contexts. For example, several papers have been carried out to survey the LLM for/with specific optimization topic (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]) and others review the general LLM applications (Hadi et al., 2024) or the application on specific problems including electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), soft engineering (Hou et al., 2023), recommendation (Wu et al., 2024b) and agent (Wang et al., 2024c). This paper aims to fill this gap by providing a systematic review with a multi-dimensional taxonomy of the current state of LLMs in algorithm design. We will explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs in enhancing algorithm design and lays the groundwork for further innovations in this exciting field. We expect it to be a helpful resource for both new researchers entering the field and experienced experts looking for a consolidated and systematic update on current developments. The contributions of this paper are as follows:
  • Systematic Review of LLM4AD: We present the first systematic review of the developments in using LLMs for algorithm design, covering a significant corpus of 180+ highly related research papers published in the last three years.
@@
  • Development of a Multi-dimensional Taxonomy: We introduce a multi-dimensional taxonomy that categorizes the works and functionalities of LLM4AD into four distinct dimensions: 1) Roles of LLMs in algorithm design, which delineates how these models contribute to or enhance algorithm design; 2) Search methods, which explores the various approaches used by LLMs to navigate and optimize search spaces in algorithm design; 3) Prompt methods, which examines how diverse prompting strategies are used; and 4) Application domains, which identifies the key fields and industries where LLMs are being applied to solve complex algorithmic challenges. This taxonomy not only clarifies the landscape but also aids in identifying gaps and opportunities for future research.
@@
  • Discussion on Challenges and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. Furthermore, we propose potential future research directions, including developing domain-specific LLMs, exploring multi-modal LLMs, facilitating human-LLM interaction, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automatic algorithm design, and benchmarking for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
@@ 2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We do not intend to cover all the literature on LLMs and algorithms. We delineate the scope of our survey as follows:
  • The term Large Language Models refers to language models of sufficient scale. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023). 1) Studies employing smaller models for algorithm design, such as those in conventional model-based and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. While it is challenging to precisely define “large” models, most leading-edge LLMs contain over one billion parameters (Zhao et al., 2023; Minaee et al., 2024). 2) Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
@@
  • Stage I Data Extraction and Collection: We collect the related papers through Google Scholar, Web of Science, and Scopus. The logic of our search is the title must include any combinations of at least one of the following two groups of words “LLM”, “LLMs”, “Large Language Model”, “Large Language Models” and “Algorithm”, “Heuristic”, “Search”, “Optimization”, “Optimizer”, “Design”, “Function” (e.g., LLM and optimization, LLMs and algorithm). After removing duplicated papers, we ended up with 850 papers as of July 1, 2024.
  • Stage II Abstract Scanning: We check the title and abstract of each paper to efficiently exclude irrelevant papers. The criteria used for exclusion include these papers that are not in English, not for algorithm design and not using large language models. After scanning, 260 papers are remaining.
  • Stage III Full Scanning: We thoroughly review each manuscript to exclude papers lacking relevant content. After scanning, there are 160 papers left.
  • Stage IV Supplementation: We append some related works manually according to our past knowledge in this field to avoid missing any important contributions. After integrating the additional papers, we got 180+ papers in the end.
@@
Figure 3. The word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It features the top 80 keywords, organized into four color-coded clusters.
We will first present an overview of the LLM4AD paper list and then present a taxonomy to systematically review the progress. In addition to the organized list of papers, we also incorporate some important publications released after July 1, 2024.
@@ 2.2. Overview
Fig. 2(a) illustrates the trend in the number of papers published over time, with the timeline expressed in months. The graph shows a marked rise in research activity related to LLM4AD, particularly noting that most of the studies have been conducted in the last year. This suggests that LLM4AD is an emerging field, and we expect a significant increase in research output in the near future as scholars from diverse fields become aware of its considerable potential.
@@ 2.3. Taxonomy
This paper presents a taxonomy organized into four dimensions: 1) LLM Roles, 2) Search Methods, 3) Prompt Methods, and 4) Applications.
@@
  • Search Methods: This category involves the search methods used including the basic sampling methods and more complex approaches like evolutionary algorithms, and uncertainty-guided methods.
@@ 3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an algorithm framework to generate and refine solutions (Fig. 4). The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions and good flexibility in prompt engineering (Yang et al., 2024; Liu et al., 2023a, 2024c), which might be infeasible for traditional optimizers. However, due to their black-box nature, they usually lack interpretability and face challenges in large-scale generation (Liu et al., 2023a).
@@
One of the initial efforts to utilize LLMs as optimizers in algorithm design is by Yang et al. (2024). They leverage the in-context learning capabilities of LLMs to generate novel solutions for specific problems based on previously evaluated solutions. This method is applied iteratively to refine solutions further. Yang et al. (2024) have successfully demonstrated this technique across various domains, including continuous and combinatorial optimization, as well as machine learning tasks.
@@
From an evolutionary algorithm (EA) perspective, using LLMs to generate solutions from existing data can be seen as analogous to search operators in EA. For instance, Liu et al. (2023a) introduce the use of LLMs as evolutionary operators to tackle multi-objective problems. This method involves breaking down a multi-objective problem into simpler single-objective tasks, with LLMs acting as black-box search operators for each sub-problem to suggest new solutions. In a related study, Liu et al. (2024c) explore the integration of LLMs within EAs, not just for generating solutions but also for guiding selection, crossover, and mutation processes. Meanwhile, Brahmachary et al. (2024) propose a new population-based evolutionary framework that includes both exploration and exploitation pools, with solutions being exchanged during the optimization process and LLMs generating solutions for both pools.
Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their EvoLLM method, which shows superior performance over baseline algorithms in synthetic black-box optimization functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms could further improve the performance of EvoLLM. Moreover, Custode et al. (2024) present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
@@
Beyond traditional optimization tasks, LLMaO has been widely adopted in prompt engineering for LLMs, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models (typically LLMs). Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity (Wang et al., 2023b), and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors (Guo et al., 2023), have been explored. Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. To address these issues, they proposed a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models.
@@ 3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models (Fig. 5) to predict the outcomes or responses of solutions, functioning either in a classification or regression context (Hao et al., 2024). Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses based on the training they have received on vast datasets. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies in the data and complicated representations (Egami et al., 2024). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
@@
The majority of LLMaP works use LLMs as pre-trained models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024e) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023) introduce CodonBERT for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and was trained on over 10 million mRNA sequences from various organisms. Soares et al. (2024) demonstrate the use of transformer-based models in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a).
@@
For classification, Hao et al. (2024) introduce LAEA, which employs LLMs as surrogate models within evolutionary algorithms for both regression and classification, eliminating the need for costly model training. In another study, Chen et al. (2023b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance. Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in Reinforcement Learning (RL). Their method demonstrated effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.